Ten more than half the babies born at Mountainside Hospital last week were girls. There we 16 boys born. How many babies were born at the Hospital last week.
Answer by drk(1908)

We can translate this sentence
"Ten more than half the babies born at Mountainside Hospital last week were girls"
as follows
(i) G = (1/2)T + 10
where G = girls and T = total
(ii) B = 16.
(iii)T = G + B
--
substitute (ii) into (iii) and we get
(iv) T = G + 16
substitute (Iv) into (i) and we get
G = (1/2)(G+16) + 10
or
G = (1/2)G + 8 + 10
which is
(1/2)G = 18
So, G = 36.
girls = 36, boys = 16, total = 52